# SMaRTT-REPS: Sender-based Marked Rapidly-adapting Trimmed & Timed Transport with Recycled Entropies

### v1

@@ 3.2.5. Overall Picture
Algorithm 1 summarizes the main SMaRTT logic. So far, we discussed lines 19-27, while we discuss the remaining parts in the next sections. It is worth remarking that SMaRTT has a low memory footprint (19 bytes per flow, and 28 bytes globally) to ensure scalability and ease of hardware implementation on NICs.
@@ 3.3. QuickAdapt
The four cases presented in Sec. 3.2 describe how SMaRTT reacts when receiving an ACK. However, if packets are dropped, SMaRTT reacts to such heavy congestion by quickly adjusting the congestion window using a novel technique, which we call QuickAdapt. QuickAdapt behaves similarly to receiver-based CC algorithms without needing special support on the receiver side (e.g., pull queues) or per-packet states like in BBR (Cardwell et al., 2017).
By default, SMaRTT relies on packet trimming to detect packet drops. Traditionally, when using packet trimming, the switch removes the payload and forwards the header if the queue size exceeds a threshold (Handley et al., 2017). This allows congestion information to be communicated to the receiver even on congested networks. With SMaRTT, we trim packets only if they would otherwise be dropped because of a full buffer.
QuickAdapt counts how many bytes have been received over the last and, if a loss is detected, it sets the congestion window to that value. It is possible to scale that value by a constant () depending on the desired queue occupancy (see Sec. 3.5). Because the sudden change caused by QuickAdapt already factors in the subsequent feedback, the receiver ignores the congestion information carried by all the unacked packets received from when QuickAdapt is triggered. By doing so, QuickAdapt does not overreact to congestion and, for the same reason, it is applied at most once per target RTT. The QuickAdapt pseudocode is presented in Algorithm 2, while the interaction with the rest of the SMaRTT logic is visible in the previously presented Algorithm 1. Figure 4 shows a simplified example of QuickAdapt in action with a 2:1 incast.
@@
Although packet trimming can be implemented on many existing switches (Adrian et al., 2022), we also want SMaRTT to be compatible with switches lacking such support. If trimming is not supported, SMaRTT falls back to a simple timeout-based approach, requiring minimal changes to the overall logic (it reacts to timeouts like it does on the reception of a trimmed packet). We compare SMaRTT performance with and without trimming in Sec. 4.2. Because SMaRTT strives to keep the RTT below the target one, it uses aggressive timeouts to detect packet losses. When using timeouts instead of trimming, we also double the value of the constant. By doing so, SMaRTT reacts more aggressively to congestion and reduces the probability of packet drops, which can decrease performance and reduce link efficiency.
@@ 3.4. FastIncrease
FastIncrease quickly reclaims available bandwidth after the termination of some flows (e.g., in case of incast with uneven sizes). FastIncrease is triggered when at least subsequent bytes did not experience any congestion (i.e., the RTT is close to the base RTT, and the ACKs are not ECN marked). In that case, as long as FastIncrease stays active, the congestion window is increased by MTUs for each ACK. is a constant that we set equal to (after proper tuning) (see Sec. 3.5). We show the FastIncrease pseudocode in Algorithm 3 and how it interacts with the rest of the SMaRTT logic in Algorithm 1.
@@ 3.5. Parameter Selection
SMaRTT enforces the congestion window size within a predefined range. We assume a maximum window size of 1.25 BDP (higher than 1 to manage transient bursts) and a minimum window size of 1 MTU. Modern link speeds and delays (with a BDP around 900KiB (Hoefler et al., 2023) and 4KiB MTUs) allow incast with hundreds (200 in this example) of senders targeting a single receiver to be easily supported. Higher incast degrees can be gracefully handled by SMaRTT at reduced link efficiency due to the potentially higher number of trimmed or dropped packets. Alternatively, SMaRTT can integrate a pacer-based solution similar to the one introduced by Swift for cases when the window would drop below one MTU (Kumar et al., 2020). However, it is worth remarking that such high-degree incast are unlikely to appear in modern AI workloads where traffic is dominated by highly optimized collective operations that limit or avoid incast (Hoefler et al., 2022).
@@
We test SMaRTT using RED queueing discipline, with and set to and of the switch queue size respectively (Floyd and Jacobson, 1993).
@@
is selected to be equal to the base RTT. Because we assume a queue size equal to 1 BDP, this corresponds to having half-full switch queues, thus halfway between and . By doing so, SMaRTT allows the REPS load balancing to react by forwarding packets on different paths, before taking more drastic actions involving a reduction of the congestion window.
@@
3.6. Recycled Entropy Packet Spray (REPS)
We couple SMaRTT with a novel per-packet load balancing mechanism, which we call REPS, which does not need any switch support besides ECMP (Hopps, 2009). With ECMP, each switch computes a hash function over some packet header fields to determine which path to forward the packet on. One of these fields is the entropy (e.g., the source port or the IPv6 Flow Label), which REPS changes per-packet.
For the first of packets, REPS explores all possible entropies. The receiver copies the entropies from the received packet to the ACK, thus forwarding them back to the sender. Whenever an ACK arrives, if the ACK is not ECN-marked, the entropy carried by the ACK is cached. Intuitively, because the corresponding data packet did not experience congestion, we can send other packets on the same path, and REPS will thus reuse it for subsequent packets. Otherwise, if the ACK was ECN-marked, REPS selects the next entropy. This can be further extended by defining additional policies (e.g., excluding bad entropies after some time or storing entropies in a priority queue based on their RTTs and/or ECN marking), but we do not explore this further in this work. We evaluate REPS’ advantages versus traditional packet spraying and ECMP in Sec. 4.1. Algorithm 4 describes the REPS pseudocode, and Fig. 5(a) shows an example of REPS’ behavior.
@@ 3.6.1. Interaction between CC and load balancing
REPS can reduce the impact of fabric congestion (e.g., caused by ECMP hashing collisions) without reducing the congestion window size. This is especially relevant in permutation workloads running on non-oversubscribed networks where, in principle, the congestion can be avoided by properly balancing the traffic. To allow REPS to react to such congestion, SMaRTT does not react to ECN-marked packets unless several marked packets are received. In a nutshell, SMaRTT keeps an exponential moving average tracking ECN-marked packets by updating, for each packet, a variable . SMaRTT does not decrease the congestion window as long as (i.e. 25 percent or more of the acks indicate congestion, which is a sufficiently large threshold to rule out congestion caused by transient load imbalance). We call this feature Wait to Decrease (WTD), and we show in Fig. 5(b) how it significantly reduces the FCT of a permutation workload running on a 1,024 nodes non-oversubscribed fat tree.
@@ 4. Evaluation
We ran simulations using and extending the htsim packet-level network simulator (Handley et al., 2017). We also performed a set of experiments (Sec. 4.1.1) on a real deployment, which we discuss later. Our simulations consider four different fat-tree topologies, two non-blocking (one with 1,024 and the other with 128 nodes), and three with 1,024 nodes and 2:1, 4:1, and 8:1 oversubscription ratios, due to the relevance of such topologies in production datacenters (Singh et al., 2015). We set a 4KiB MTU size, a 800 Gbps network bandwidth, and a 400ns switch traversal latency, consistent with the latest generation of switches (Broadcom, 2024b; De Sensi et al., 2020). For the sake of simplicity, we assume that all links have the same length and, hence, the same latency, equal to 600ns.
@@
We evaluate SMaRTT-REPS against EQDS, Swift, MPRDMA, and BBR. To have a fair comparison and focus on congestion control, we enable trimming and multi-pathing for all the algorithms. Indeed, running Swift or BBR in their vanilla version using a single path would result in bad performance, mostly due to routing rather than congestion control. We assessed this experimentally by comparing the single-path vanilla version of Swift (using PLB (Qureshi et al., 2022)), a version using oblivious packet spraying, and one using REPS. We found out that the version using REPS consistently outperforms the others since PLB reacts to congestion only after several RTTs, and is thus not suitable for the scenarios we target. We thus use REPS for all the algorithms and all the experiments, except when stated otherwise. We do not compare SMaRTT with DCTCP and DCQCN since it has already been shown that EQDS consistently outperforms them (Olteanu et al., 2022).
@@
In the following, we first analyze the performance gain obtained through the use of REPS (Sec. 4.1), as well as the advantages of using packet trimming (Sec. 4.2). Then, we compare SMaRTT’s performance with state-of-the-art algorithms on incast (Sec. 4.3), permutation (Sec. 4.4), and alltoall (Sec. 4.5) patterns.
4.1. REPS
In Fig. 6(a), we compare SMaRTT’s performance using different load balancing algorithms (REPS, ECMP (Hopps, 2009), PLB (Qureshi et al., 2022; Kabbani et al., 2014), and oblivious packet spraying (Dixit et al., 2013)). We simulate a 32MiB permutation pattern on the 4:1 oversubscribed fat tree. We observe that, due to better fairness, REPS improves the completion time by roughly 10% compared to oblivious spraying. We also observe that with per-flow ECMP, the job completion time is 1.5x times higher than that of packet spraying and REPS due to more flow collisions (not shown in the plot). PLB also struggle but it is able to achieve a much better fairness compared to ECMP. Moreover, by analyzing the packet RTT distribution in Fig. 6(b). We observe that REPS is characterized by RTTs closer to the target one and with a lower variability, indicating better network utilization.
4.1.1. Handling Faults and Network Asymmetries
REPS is also effective in dealing with network faults and asymmetries. To assess such capabilities, we implemented REPS in a production-grade FPGA-based RDMA-capable NIC. We then compared REPS against a baseline oblivious packet spray approach. The experimental setup consists of a two-tier fat-tree network with 100G FPGA-based NICs and Tomahawk3-based switches. Tier 0 (T0) downlinks are 100G and baseline Tier0-Tier1 uplinks are 400G. Endpoints run 4MB ring-based AllReduce operations on RDMA, with the logical ring laid out such that all connections traverse the Tier 1 (T1) spine and maximize pressure on the network.
In the first experiment, we connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptiveness of REPS, we changed the link speed of one T0-T1 link from 400G to 200G, creating asymmetry in the network. Fig. 7(a) shows the per-flow goodput as observed by the application. Oblivious packet spray sends packets across all paths (including those crossing the 200G link) with equal probability and is ultimately capped by the slower 200G path. Indeed, the congestion signal produced from the slower path traffic limits the total flow goodput to the bottlenecked path (thus severely underutilizing the remaining capacity of the 400G links).
In contrast, REPS can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, REPS can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target. We also report in Fig. 7(b) hardware-collected FCT distributions, showing that REPS can maintain a lower and tighter FCT distribution, significantly reducing average and tail FCTs.
To demonstrate the robustness and resilience of REPS in the context of link failures, Fig. 7(c) shows total packet drops (average across four runs each) observed in a large-scale 128-endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), the oblivious spraying continues sending packets across all paths (including those affected by the link that went down). On the other hand, REPS quickly adapts (within the order of an RTT) and avoids sending packets down the affected paths as the entropy cache is replenished from packets traversing healthy paths. In these sample runs, REPS exhibited a much more predictable behavior and reduced the impact of link failure, providing an order of magnitude reduction in the total number of packet drops.
@@ 4.1.1. Handling Faults and Network Asymmetries
As discussed in Sec. 3.3, SMaRTT relies on timeouts to detect packet drops when the switches do not support trimming. We now evaluate the impact of the lack of trimming. In Fig. 8, we report an incast scenario on small (512 KiB) messages, showing that the lack of trimming increases the FCT by slightly more than one base RTT (equal to 8.6 s in our setup).
@@ 4.3. Incast
Fig. 10 shows the relative performance of SMaRTT in incast scenarios ranging from an 8:1 incast to a 100:1 incast for varying message sizes. We note that EQDS, being a receiver-based mechanism, can easily achieve good performance. Indeed, it can precisely communicate the window size to each sender and schedule them to be perfectly fair.
On the other hand, sender-based mechanisms like MPRDMA, SMaRTT, and Swift perform slightly worse, particularly true for medium sized messages. Small messages do not overrun switch buffers, whereas for large enough messages all the algorithms have enough time to converge to the correct rate. The zone in between, usually for messages slightly smaller than BDP, is subject to random drop events that affect sender-based mechanisms more. Moreover, MPRDMA seems to have slightly worse performance due to its lack of fairness features. Finally, BBR struggles for medium-sized messages since it needs more time to converge to the correct rate.
@@ 4.4. Permutation
In Fig. 1, we show the results for a 2 MiB and 32 MiB permutation traffic pattern on an 8:1 oversubscribed fat tree with 1,024 nodes. It is worth noting that, when using a relatively small message of 2 MiB, just slightly bigger than the BDP, SMaRTT outperforms all other algorithms while being fair (it can quickly react thanks to ECN). In Fig. 11(a) and Fig. 11(b), we show what happens on 2:1 and 4:1 oversubscribed fat trees respectively. In these cases, SMaRTT still outperforms the other algorithms, but not as much as in other cases, a smaller oversubscription ratio is easier to manage even for simpler CC algorithm.
In Fig. 11(c) we report instead the results for a scenario where multiple permutations run in parallel (which can be the case with a Butterfly AllReduce or a windowed AllToAll). In this case, the performance gap between SMaRTT and EQDS increases up to 50%, while all the other sender based-algorithms perform better but are still about 10% slower than SMaRTT. Finally, we show in Fig. 11(d) a scenario with 2MiB flows, except for one that is bigger and sends 4MiB. This scenario shows how SMaRTT outperforms other sender-based mechanisms by almost 30% thanks to its capacity to reclaim bandwidth quickly (see Sec. 3.4). This is also the case for EQDS, which outperforms MPRDMA, BBR, and Swift since it can also reclaim unused bandwidth quickly.
Generally, we can notice that for high oversubscription ratios, the performance of EQDS degrades even more due to the high number of trimmed packets. To further emphasize this point, we measure the number of trimmed packets for EQDS and report that, in the worst case scenario just mentioned, it can generate up to 155x more trimmed packets than SMaRTT, resulting in wasted bandwidth and resources. The EQDS authors state that in such cases a CC algorithm on the sender side would benefit its performance, and we explore that in Sec. 5.1.
@@ 4.5. Alltoall
We also evaluate performance for alltoall communication, a collective operation commonly used in AI workloads (Hoefler et al., 2022). Due to the large number of messages generated, we use the smaller topology to reduce simulation times. We use a windowed algorithm for the alltoall to have at least active flows per node at any time. We show the results in an oversubscribed fat tree, as a non blocking tree topology can theoretically handle alltoall traffic without problems for congestion control. We report the results of our evaluation in Fig. 12. We observe that EQDS performance drops as the number of parallel connection increases, consistent with the multiple permutation results in Fig. 11(c). On the other hand, sender-based mechanisms handle that case better, with SMaRTT taking the lead being only 6% behind the ideal time.
@@ 5. Discussion
5.1. Augmenting EQDS with SMaRTT
One of the main drawbacks of using a receiver-based mechanism like EQDS, as shown in the previous results, is its bad management of fabric congestion. As suggested in the EQDS paper, it should be possible to improve EQDS’ performance by complementing it with a sender-based CC algorithm. To do so, we explore using EQDS with SMaRTT, as well as with a DCTCP-style CC algorithm (MPRDMA). While running in combination with EQDS, the role of the sender-based congestion control is to limit its sending rate by capping the congestion window and adapting as needed.
In theory, this combination should give us most of the benefits of both world: perfect incast management thanks to the receiver-based CC and the ability to deal with the other forms of congestion thanks to a solid sender-based CC. In this work, we only show a simple scenario where such combinations seem beneficial for EQDS. A more in-depth analysis of such a combination is necessary but outside the scope of this paper.
In Fig. 13 we report a permutation scenario on an 8:1 oversubscribed fat tree, with 2 MiB and 32 MiB flows. We can see that SMaRTT improves EQDS performance. For 2MiB flows, the improvement introduced by SMaRTT is higher than that of MPRDMA, thanks to its better fairness. When running with 32 MiB flows except for one being 64 MiB, we can also see how we do well on our own but also help EQDS.

### v3

@@ 3.2.5. Overall Picture
Algorithm 1 summarizes the main FASTFLOW logic. So far, we discussed lines 19-27, while we discuss the remaining parts in the next sections. It is worth remarking that FASTFLOW has a low memory footprint (19 bytes per flow, and 28 bytes globally) to ensure scalability and ease of hardware implementation on NICs.
@@ 3.3. QuickAdapt
The four cases presented in Sec. 3.2 describe how FASTFLOW reacts when receiving an ACK. However, if packets are dropped, FASTFLOW reacts to such heavy congestion by quickly adjusting the congestion window using a novel technique, which we call QuickAdapt. QuickAdapt behaves similarly to receiver-based CC algorithms without needing special support on the receiver side (e.g., pull queues) or per-packet states like in BBR (Cardwell et al., 2017).
By default, FASTFLOW relies on packet trimming to detect packet drops. Traditionally, when using packet trimming, the switch removes the payload and forwards the header if the queue size exceeds a threshold (Handley et al., 2017). This allows congestion information to be communicated to the receiver even on congested networks. With FASTFLOW, we trim packets only if they would otherwise be dropped because of a full buffer.
QuickAdapt counts how many bytes have been received over the last and, if a loss is detected, it sets the congestion window to that value. It is possible to scale that value by a constant () depending on the desired queue occupancy (see Sec. 3.5). Because the sudden change caused by QuickAdapt already factors in the subsequent feedback, the receiver ignores the congestion information carried by all the unacked packets received from when QuickAdapt is triggered. By doing so, QuickAdapt does not overreact to congestion and, for the same reason, it is applied at most once per target RTT. The QuickAdapt pseudocode is presented in Algorithm 2, while the interaction with the rest of the FASTFLOW logic is visible in the previously presented Algorithm 1. Figure 4 shows a simplified example of QuickAdapt in action with a 2:1 incast.
@@
Although packet trimming can be implemented on many existing switches (Adrian et al., 2022), we also want FASTFLOW to be compatible with switches lacking such support. If trimming is not supported, FASTFLOW falls back to a simple timeout-based approach, requiring minimal changes to the overall logic (it reacts to timeouts like it does on the reception of a trimmed packet). We compare FASTFLOW performance with and without trimming in Sec. 4.1. Because FASTFLOW strives to keep the RTT below the target one, it uses aggressive timeouts to detect packet losses. When using timeouts instead of trimming, we also double the value of the constant. By doing so, FASTFLOW reacts more aggressively to congestion and reduces the probability of packet drops, which can decrease performance and reduce link efficiency.
@@ 3.4. FastIncrease
FastIncrease quickly reclaims available bandwidth after the termination of some flows (e.g., in case of incast with uneven sizes). FastIncrease is triggered when at least subsequent bytes did not experience any congestion (i.e., the RTT is close to the base RTT, and the ACKs are not ECN marked). In that case, as long as FastIncrease stays active, the congestion window is increased by MTUs for each ACK. is a constant that we set equal to (after proper tuning) (see Sec. 3.5). We show the FastIncrease pseudocode in Algorithm 3 and how it interacts with the rest of the FASTFLOW logic in Algorithm 1.
@@ 3.5. Parameter Selection
FASTFLOW enforces the congestion window size within a predefined range. We assume a maximum window size of 1.25 BDP (higher than 1 to manage transient bursts) and a minimum window size of 1 MTU. Modern link speeds and delays (with a BDP around 900KiB (Hoefler et al., 2023) and 4KiB MTUs) allow incast with hundreds (200 in this example) of senders targeting a single receiver to be easily supported. Higher incast degrees can be gracefully handled by FASTFLOW at reduced link efficiency due to the potentially higher number of trimmed or dropped packets. Alternatively, FASTFLOW can integrate a pacer-based solution similar to the one introduced by Swift for cases when the window would drop below one MTU (Kumar et al., 2020). However, it is worth remarking that such high-degree incast are unlikely to appear in modern AI workloads where traffic is dominated by highly optimized collective operations that limit or avoid incast (Hoefler et al., 2022).
@@
We test FASTFLOW using RED queueing discipline, with and set to and of the switch queue size respectively (Floyd and Jacobson, 1993).
@@
is selected to be equal to the base RTT. Because we assume a queue size equal to 1 BDP, this corresponds to having half-full switch queues, thus halfway between and . By doing so, FASTFLOW allows the loead balancer to react by forwarding packets on different paths, before taking more drastic actions involving a reduction of the congestion window.
@@ 3.5.1. Interaction between CC and load balancing
An adaptive load balancer can reduce the impact of fabric congestion (e.g., caused by ECMP hashing collisions) without reducing the congestion window size. This is especially relevant in permutation workloads running on non-oversubscribed networks where, in principle, the congestion can be avoided by properly balancing the traffic. To allow adaptive load balancer algorithms to react to such congestion, FASTFLOW does not react to ECN-marked packets unless several marked packets are received. In a nutshell, FASTFLOW keeps an exponential moving average tracking ECN-marked packets by updating, for each packet, a variable . FASTFLOW does not decrease the congestion window as long as (i.e. 25 percent or more of the acks indicate congestion, which is a sufficiently large threshold to rule out congestion caused by transient load imbalance). We call this feature Wait to Decrease (WTD).
@@ 4. Evaluation
We ran simulations using and extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider four different fat-tree topologies, two non-blocking (one with 1,024 and the other with 128 nodes), and three with 1,024 nodes and 2:1, 4:1, and 8:1 oversubscription ratios, due to the relevance of such topologies in production datacenters (Singh et al., 2015). We set a 4KiB MTU size, a 800 Gbps network bandwidth, and a 400ns switch traversal latency, consistent with the latest generation of switches (Broadcom, 2024b; De Sensi et al., 2020). For the sake of simplicity, we assume that all links have the same length and, hence, the same latency, equal to 600ns.
@@
We evaluate FASTFLOW against EQDS, Swift, MPRDMA, and BBR. To have a fair comparison and focus on congestion control, we enable trimming and multi-pathing for all the algorithms. Indeed, running Swift or BBR in their vanilla version using a single path would result in bad performance, mostly due to routing rather than congestion control.
@@
As discussed in Sec. 3.3, FASTFLOW relies on timeouts to detect packet drops when the switches do not support trimming. We now evaluate the impact of the lack of trimming. In Fig. 6, we report an incast scenario on small (512 KiB) messages, showing that the lack of trimming increases the FCT by slightly more than one base RTT (equal to 8.6 s in our setup).
@@ 4.2. Incast
Fig. 5 shows the relative performance of FASTFLOW in incast scenarios ranging from an 8:1 incast to a 100:1 incast for varying message sizes. We note that EQDS, being a receiver-based mechanism, can easily achieve good performance. Indeed, it can precisely communicate the window size to each sender and schedule them to be perfectly fair.
On the other hand, sender-based mechanisms like MPRDMA, FASTFLOW, and Swift perform slightly worse, particularly true for medium sized messages. Small messages do not overrun switch buffers, whereas for large enough messages all the algorithms have enough time to converge to the correct rate. The zone in between, usually for messages slightly smaller than BDP, is subject to random drop events that affect sender-based mechanisms more. Moreover, MPRDMA seems to have slightly worse performance due to its lack of fairness features. Finally, BBR struggles for medium-sized messages since it needs more time to converge to the correct rate.
@@ 4.3. Permutation
In Fig. 1, we show the results for a 2 MiB and 32 MiB permutation traffic pattern on an 8:1 oversubscribed fat tree with 1,024 nodes. It is worth noting that, when using a relatively small message of 2 MiB, just slightly bigger than the BDP, FASTFLOW outperforms all other algorithms while being fair (it can quickly react thanks to ECN). In Fig. 8(a) and Fig. 8(b), we show what happens on 2:1 and 4:1 oversubscribed fat trees respectively. In these cases, FASTFLOW still outperforms the other algorithms, but not as much as in other cases, a smaller oversubscription ratio is easier to manage even for simpler CC algorithm.
In Fig. 8(c) we report instead the results for a scenario where multiple permutations run in parallel (which can be the case with a Butterfly AllReduce or a windowed AllToAll). In this case, the performance gap between FASTFLOW and EQDS increases up to 50%, while all the other sender based-algorithms perform better but are still about 10% slower than FASTFLOW. Finally, we show in Fig. 8(d) a scenario with 2MiB flows, except for one that is bigger and sends 4MiB. This scenario shows how FASTFLOW outperforms other sender-based mechanisms by almost 30% thanks to its capacity to reclaim bandwidth quickly (see Sec. 3.4). This is also the case for EQDS, which outperforms MPRDMA, BBR, and Swift since it can also reclaim unused bandwidth quickly.
Generally, we can notice that for high oversubscription ratios, the performance of EQDS degrades even more due to the high number of trimmed packets. To further emphasize this point, we measure the number of trimmed packets for EQDS and report that, in the worst case scenario just mentioned, it can generate up to 155x more trimmed packets than FASTFLOW, resulting in wasted bandwidth and resources. The EQDS authors state that in such cases a CC algorithm on the sender side would benefit its performance, and we explore that in Sec. 5.1.
@@ 4.4. Alltoall
We also evaluate performance for alltoall communication, a collective operation commonly used in AI workloads (Hoefler et al., 2022). Due to the large number of messages generated, we use the smaller topology to reduce simulation times. We use a windowed algorithm for the alltoall to have at least active flows per node at any time. We show the results in an oversubscribed fat tree, as a non blocking tree topology can theoretically handle alltoall traffic without problems for congestion control. We report the results of our evaluation in Fig. 9. We observe that EQDS performance drops as the number of parallel connection increases, consistent with the multiple permutation results in Fig. 8(c). On the other hand, sender-based mechanisms handle that case better, with FASTFLOW taking the lead being only 6% behind the ideal time.
@@ 5. Discussion
5.1. Augmenting EQDS with FASTFLOW
One of the main drawbacks of using a receiver-based mechanism like EQDS, as shown in the previous results, is its bad management of fabric congestion. As suggested in the EQDS paper, it should be possible to improve EQDS’ performance by complementing it with a sender-based CC algorithm. To do so, we explore using EQDS with FASTFLOW, as well as with a DCTCP-style CC algorithm (MPRDMA). While running in combination with EQDS, the role of the sender-based congestion control is to limit its sending rate by capping the congestion window and adapting as needed.
In theory, this combination should give us the benefits of both world: perfect incast management thanks to the receiver-based CC and the ability to deal with the other forms of congestion thanks to a solid sender-based CC. In this work, we only show a simple scenario where such combinations seem beneficial for EQDS. A more in-depth analysis of such a combination is necessary but outside the scope of this paper.
In Fig. 10 we report a permutation scenario on an 8:1 oversubscribed fat tree, with 2 MiB and 32 MiB flows. We can see that FASTFLOW improves EQDS performance. For 2MiB flows, the improvement introduced by FASTFLOW is higher than that of MPRDMA, thanks to its better fairness. When running with 32 MiB flows except for one being 64 MiB, we can also see how we do very well on our own but also help EQDS. In this case having FastIncrease is key to recover quickly the bandwidth for the larger flow.